Cryo Problems 07/12

The place to discuss the LHC. Commissioning, operation, issues, events ....
Post Reply
Danny252
LHCPortal Guru
Posts: 169
Joined: Mon Nov 23, 2009 4:45 pm

Cryo Problems 07/12

Post by Danny252 » Mon Dec 07, 2009 11:42 am

At 0425 this morning, Sector 45's Cryo system indicated a problem with the warm compressor (which is..?).

At 0439, all sectors other than 23 had tripped - to quote the logbook: "Now all the sectors except S23 are tripping! Why?"

This was soon resolved to be a brief electrical fault that affected multiple systems - eventually all the Cryo systems came back apart from those at Point 4, and at 0630 when most other problems had been solved, the magnet temperatures in Point 4 were still rising. The problem is still ongoing, as far as I can tell.

Also, 24 of the 100 collimators were affected by the power glitch - the LHC COLL eLogBook gives a list of these.

User avatar
DCWhitworth
LHCPortal Guru
Posts: 584
Joined: Mon Nov 30, 2009 8:13 am
Location: Norwich, UK

Re: Cryo Problems 07/12

Post by DCWhitworth » Mon Dec 07, 2009 10:23 pm

Small Cryo issue on the inner triple magnet in sector 56, showing yellow HCC field page, was red on Cryo page 1 but looks OK now.
DC

The LHC - One ring to rule them all !

User avatar
seabass
Posts: 18
Joined: Sun Dec 06, 2009 2:36 pm
Location: Chicago

Re: Cryo Problems 07/12

Post by seabass » Mon Dec 07, 2009 11:53 pm

I know. Going on the LHC latest news reads like a best selling love/hate drama.

00:00-02:30: Four bunch setup.
a) 1.5e10 p in beam1 in 4 bunches. Fully protected machine. 25h beam lifetime. Correctly dumped.
b) At 01:32: Two beams with each 4 bunches stored. B1: 1.4e10. B2: 1.5e10.
c) Dump tests show normal loss patterns both for bunched and unbunched beams.
02:30-06:30: Injecting for stable beams. Finally successful around 6:30 after resolving a lifetime problem with beam 2. 4 on 4 circulating and colliding with reasonable lifetime for both beams.
06:55-11:00: STABLE BEAMS!



Monday 7th...



Since 00:30: Beam off.
a) Loss of cryo maintain in sector 81 led to dump of stable beams.
b) Electrical perturbation at 04:25 during recovery. Loss of warm compressor in IR4. Magnet trips. Other effects on machine elements.
c) Recovery estimated between 22:00 and 24:00.




:D

Post Reply